# A Third-Order Relativistic Algebraic Diagrammatic Construction Method for Double Ionization Potentials: Theory, Implementation, and Benchmark

### v1

@@ IV.5 The level of treatment of relativistic Hamiltonians
We now investigate the consequences of neglecting relativistic effects in DIP calculations, as well as the impact of incorporating higher-order relativistic Hamiltonians. In Table 6, we present the DIP values computed using non-relativistic, Dirac–Coulomb (DC), Dirac–Coulomb–Gaunt (DCG), and Dirac–Coulomb–Breit (DCB) Hamiltonians for selected noble gas atoms. One clear observation is that relativistic treatment is crucial for correctly capturing spin–orbit splittings. Notably, the $^{3}P_{2}$ state of Rn is overestimated by about 1.4 eV in the non-relativistic framework, whereas the DC Hamiltonian brings the error down to just 0.43 eV. A similar trend is seen in the doubly ionized states of Xe. Incorporating the Gaunt correction into the DC Hamiltonian leads to a slight reduction in DIP values across all considered states, with the largest decrease of 0.04 eV observed for the $^{1}S_{0}$ state of the Xe atom. The Breit correction produces a similar trend, also causing a maximum decrease of 0.04 eV. The same set of calculations is then applied to the diatomic molecules considered in this study. The corresponding results are presented in Table 7. The inclusion of relativistic effects using the DC Hamiltonian reduces the MAE from 0.15 eV to 0.11 eV. The importance of incorporating relativistic effects is more pronounced for molecules containing heavy elements. When the Gaunt correction is added to the DC Hamiltonian, the DIP values either remain unchanged or show a slight decrease, with a maximum shift of just 0.02 eV. A similar trend is observed upon inclusion of the Breit correction. The statistical error parameters for both DCG and DCB results remain nearly identical to those of the DC calculations, indicating that the effects of Gaunt and Breit corrections have minimal impact in this context.
V Conclusion
@@
In summary, we have presented the theory, implementation, and benchmarking of a third-order relativistic ADC method for calculating double ionization potentials. The extension to the third order significantly increases the accuracy over the DIP-ADC(2) method. The computational cost of the relativistic DIP-ADC(3) method can be reduced by using a two-component X2CAMF Hamiltonian combined with the FNS framework and CD-based treatment of the two-electron integrals. The X2CAMF-DIP-ADC(3) method gives almost identical DIP values as those obtained with a four-component DC Hamiltonian. The use of the FNS approximation allows one to truncate the virtual space for the DIP-ADC calculation with systematic controllable accuracy. Additionally, the use of CD and FNS approximation allowed us to employ large basis sets like dyall.av4z, even for heavy elements. Our analysis shows that lower-order approximations like ADC(2) are not sufficient for reliable predictions of DIP due to their inadequate treatment of the $3h1p$ sector. The DIPs computed using the DC Hamiltonian at the ADC(3) level of theory showed a MAE of just 0.11 eV when compared with the experimental results, highlighting the robustness and precision of the relativistic ADC(3) method. The agreement with the experimental results is comparable, often better than the relativistic DIP-EOM-CCSD($3h1p$) method. The inclusion of Gaunt and Breit corrections has a minimal effect on the valence DIP values. To obtain a quantitative agreement for DIP values, one needs to include the $4h2p$ blocks in the calculation, which essentially means going to the ADC(4)Schirmer, Cederbaum, and Walter (1983); Leitner, Dempwolff, and Dreuw (2022) level. Work is in progress towards developing a lower scaling relativistic DIP-ADC(4) method.
VI SUPPLEMENTARY MATERIAL
@@
VII Acknowledgements
The authors gratefully acknowledge financial support from IIT Bombay, including the IIT Bombay Seed Grant (Project No. R.D./0517-IRCCSH0-040), CRG (Project No. CRG/2022/005672), and MATRICS (Project No. MTR/2021/000420) projects of DST-SERB; CSIR-India (Project No. 01(3035)/21/EMR-II); DST-Inspire Faculty Fellowship (Project No. DST/INSPIRE/04/2017/001730); and ISRO (Project No. RD/0122-ISROC00-004). The authors also acknowledge the IIT Bombay supercomputing facility and C-DAC resources (Param Smriti, Param Bramha, and Param Rudra) for providing computational time. S.M. acknowledges support from the Prime Minister’s Research Fellowship (PMRF).

### v2

@@ IV.5 Different levels of relativistic hamiltonian treatment
We now investigate the consequences of neglecting relativistic effects in DIP calculations, as well as the impact of incorporating higher-order terms in the relativistic Hamiltonians. In Table 6, we present the DIP-ADC(3) values computed using non-relativistic, Dirac–Coulomb (DC), Dirac–Coulomb–Gaunt (DCG), and Dirac–Coulomb–Breit (DCB) Hamiltonians for selected noble gas atoms. The Gaunt and the Breit interactions have only been considered in the Hartree-Fock step and have not been included in the subsequent integral transformation. One clear observation is that relativistic treatment is crucial for correctly describing the DIP values of heavier elements. Notably, the $^{3}P_{2}$ state of Rn is overestimated by about 1.4 eV in the non-relativistic framework, whereas the DC Hamiltonian brings the error down to just 0.43 eV. A similar trend is seen in the doubly ionized states of Xe. Incorporating the Gaunt correction into the DC Hamiltonian leads to a slight reduction in DIP values across all considered states, with the largest decrease of 0.04 eV observed for the $^{1}S_{0}$ state of the Xe atom. The change in the DIP value due to the inclusion of Breit correction is negligible over that obtained with the DCG Hamiltonian. We have also investigated the effect of the treatment of relativity for the diatomic molecules considered in this study. The corresponding results are presented in Table 7. The inclusion of relativistic effects using the DC Hamiltonian reduces the MAE from 0.15 eV in a non-relativistic calculation to 0.11 eV. The importance of incorporating relativistic effects is more pronounced for molecules containing heavy elements. When the Gaunt correction is added to the DC Hamiltonian, the DIP values either remain unchanged or show a slight decrease, with a maximum shift of just 0.02 eV. The inclusion of the Breit correction produces only marginal changes relative to the DCG Hamiltonian. The statistical error parameters for both DCG and DCB results remain nearly identical to those obtained with the DC Hamiltonian, indicating that the effects of Gaunt and Breit corrections have minimal impact in this context.
V Computational efficiency
A central aim of using CD and X2CAMF approximations is to lower the computational cost of DIP-ADC(3) calculations for large basis sets. As a test case, we calculated the DIP of the HI molecule with the dyall.av3z basis using full four-component- (4c-), FNS-4c-, and FNS-CD-X2CAMF- DIP-ADC(3) methods and compared the timings across different stages of the computation. The calculations were performed
Figure 4: Comparison of computational timings for different steps in the DIP calculation of HI using canonical 4c, FNS-4c, and FNS-CD-X2CAMF based DIP-ADC(3) methods.
using a single thread on a workstation comprising two Intel(R) Xeon(R) Silver 4216 CPUs with a clock speed of 2.10 GHz, and 512 GB of available RAM. All calculations were performed using the frozen-core approximation. For the FNS-based computations, an FNS threshold of $10^{-4.5}$ was used, while a CD threshold of $10^{-5}$ was applied in the FNS-CD-X2CAMF approach. In the full 4c calculation, 382 virtual spinors were correlated, while the FNS procedure retained only 28% of this space, corresponding to 104 virtual spinors. From Figure 4, it is clear that the integral transformation is the most time-consuming step in both the full 4c and FNS-4c calculations. In the full 4c case, this step required 8 hours, 12 minutes, and 33 seconds, whereas employing the FNS basis reduced the time to 1 hour, 5 minutes, and 16 seconds. The FNS-CD-X2CAMF approach provides further advantages by eliminating the need to generate and store three- and four-virtual integrals, yielding substantial savings in both storage and computational effort. Considering the overall timings, the full 4c method required 9 hours 12 minutes, the FNS-4c approach took about 1 hour 11 minutes, while the FNS-CD-X2CAMF scheme completed the calculation in just 2 minutes. For this particular case, the FNS-CD-X2CAMF method reproduced the full 4c DIP value up to the fourth decimal place in eV, highlighting its outstanding efficiency and its ability to overcome storage limitations.
VI Conclusion
@@
In summary, we have presented the theory, implementation, and benchmarking of a third-order relativistic ADC method for calculating double ionization potentials. The extension to the third order significantly increases the accuracy over the DIP-ADC(2) method. The computational cost of the relativistic DIP-ADC(3) method can be reduced by using a two-component X2CAMF Hamiltonian combined with the FNS framework and CD-based treatment of the two-electron integrals. The X2CAMF-DIP-ADC(3) method gives almost identical DIP values as those obtained with a four-component DC Hamiltonian. The use of the FNS approximation allows one to truncate the virtual space for the DIP-ADC calculation with systematic controllable accuracy. Additionally, the use of CD allowed us to employ large basis sets, even for heavy elements. Our analysis shows that lower-order approximations like ADC(2) are not sufficient for reliable predictions of DIP due to their inadequate treatment of the $3h1p$ configuration. The DIPs computed using the DC Hamiltonian at the ADC(3) level of theory showed a MAE of just 0.11 eV when compared with the experimental results, highlighting the robustness and precision of the relativistic ADC(3) method. The results show agreement with experiment that is comparable to, and often better than, the relativistic DIP-EOM-CCSD($3h1p$) method. The inclusion of Gaunt and Breit corrections has a minimal effect on the valence DIP values. To obtain a quantitative agreement for DIP values, one needs to include the $4h2p$ blocks in the calculation, which essentially means going to the ADC(4)Schirmer, Cederbaum, and Walter (1983); Leitner, Dempwolff, and Dreuw (2022) level of theory. Work is in progress towards developing a lower scaling relativistic DIP-ADC(4) method.
SUPPLEMENTARY MATERIAL
@@
Acknowledgements
The authors gratefully acknowledge financial support from IIT Bombay, including the IIT Bombay Seed Grant (Project No. R.D./0517-IRCCSH0-040), CRG (Project No. CRG/2022/005672), and MATRICS (Project No. MTR/2021/000420) projects of DST-SERB; CSIR-India (Project No. 01(3035)/21/EMR-II); DST-Inspire Faculty Fellowship (Project No. DST/INSPIRE/04/2017/001730); and ISRO (Project No. RD/0122-ISROC00-004). The authors also acknowledge the IIT Bombay supercomputing facility and C-DAC resources (Param Smriti, Param Bramha, and Param Rudra) for providing computational time. S.M. gratefully acknowledges support from the Prime Minister’s Research Fellowship (PMRF).